### v1

# Throughput-Optimal Scheduling Algorithms for LLM Inference and AI Agents

Abstract
As demand for Large Language Models (LLMs) and AI agents rapidly grows, optimizing systems for efficient LLM inference becomes critical. While significant efforts have targeted system-level engineering, little is explored through a mathematical modeling and queuing perspective.
In this paper, we aim to develop the queuing fundamentals for LLM inference, bridging the gap between queuing and LLM system communities. In particular, we study the throughput aspect of LLM inference systems. We prove that a large class of ’work-conserving’ scheduling algorithms can achieve maximum throughput for both individual requests and AI-agent workloads, highlighting ’work-conserving’ as a key design principle in practice. Evaluations of real-world systems show that Orca and Sarathi-serve are throughput-optimal, reassuring practitioners, while FastTransformer and vanilla vLLM are not maximally stable and should be used with caution. Our results highlight the substantial benefits queuing community can offer in analyzing and improving LLM inference systems and call for more interdisciplinary developments.
1 Introduction
Large Language Models (LLMs) have become the backbone of many AI-driven applications, requiring efficient inference serving engines to meet latency and throughput requirements. LLM inference or serving servers process diverse workloads that include prefill (prompt processing) and decode (token generation) tasks, each with distinct computational and memory bandwidth characteristics. Since these models produce tokens autoregressively [^1] A token in an autoregressive transformer is the basic input or output unit—typically a word, subword, or character—that the model sequentially predicts based on previously generated tokens., processing a single inference task requires running the model multiple times, with each iteration generating one output token. As LLMs scale in size and complexity, optimizing GPU utilization while ensuring fast response times remains a critical challenge.
To maximize throughput while meeting latency requirements, several inference systems with different scheduling algorithms have been developed. FasterTransformer (NVIDIA, 2024) introduces a request-level batch scheduler, thereby increasing throughput. Subsequently, Orca (Yu et al., 2022) and vanilla vLLM (Kwon et al., 2023) were developed to enable batching at the token level, with prioritization of prefill tokens for newly arriving requests—allowing these requests to enter the decoding stage sooner. The two approaches, however, differ in their batch composition strategies: vLLM enforces homogeneous batches containing either only prefill or only decode tokens, while Orca allows hybrid batches that mix both types. Despite these differences, both frameworks aim to maximize batch size within memory and latency constraints to improve goodput.
Non-mixed batching policies, such as those in FasterTransformer and vanilla vLLM (NVIDIA, 2024; Kwon et al., 2023), lead to suboptimal throughput due to idling GPUs when one type of token is underrepresented. Mixed batching strategies, like those in Orca (Yu et al., 2022), address this inefficiency by allowing prefill and decode tokens to coexist within the same batch, improving utilization in memory bandwidth and compute. More recently, Sarathi-Serve introduces chunk-prefill to limit the number of tokens in a single batch to avoid decode stall due to long prefill interference (Agrawal et al., 2023).
While significant advances have been made from the system community, the field lacks a cohesive theoretical framework for evaluating and comparing inference systems with different scheduling algorithms. Consequently, practitioners are often left to decide whether to switch from one system to another based on case studies on comprehensive workload benchmarking (Databricks, 2023), sometimes using profiling tools (Shen et al., 2022; NVIDIA Corporation, 2023) to understand which algorithm falls short. This is often costly to evaluate across various models, hardware setups, workload generation patterns, datasets, etc. Meanwhile, the operations research and queuing theory communities have established robust theoretical foundations and practical methodologies for scheduling and resource allocation problems, successfully implementing these across diverse domains including cloud computing (Mitzenmacher, 2001; Grosof et al., 2023; Tirmazi et al., 2020; Patke et al., 2024; Ghodsi et al., 2011; Zhang et al., 2023), operating systems and networking (Yu et al., 2021; Stoica et al., 1998; Iyer et al., 2023; Li et al., 2024), telecommunications (Whitt, 1993; Maglaris et al., 1988; Canetti & Irani, 1995), manufacturing and healthcare operations (Hu et al., 2022; Chen et al., 2025; Zychlinski et al., 2023), etc.
Motivated by this gap, we analyze LLM inference systems through queuing theory to address the fundamental question
What is the fundamental limit of the maximal throughput rate for an LLM inference system, and what classes of scheduling algorithms can achieve this limit?
We develop an analytically tractable model of LLM inference and identify a broad class of work-conserving scheduling policies that achieve maximal throughput, based on realistic system profiling and performance modeling. Our analysis confirms that systems like Sarathi-Serve and Orca are throughput-optimal, thereby offering practical assurance to system designers. In contrast, non-mixed batching approaches—such as those in FastTransformer and vanilla vLLM—are inherently suboptimal in throughput and can become unstable under moderate load. These findings contribute to a deeper theoretical foundation for LLM system design and provide a principled basis for evaluating and comparing scheduling strategies.
Our work extends batch queuing theory by addressing LLM-specific challenges: dual-phase processing with distinct resource profiles, dynamic batch formation, and request interdependencies. We further analyze AI-agent workloads, in which a network of LLM instances collaboratively processes agent-level tasks. As AI agents become increasingly prevalent and embedded in human workflows, understanding the performance characteristics of such distributed inference systems becomes critical. Our analysis identifies several regimes in which optimal throughput can still be achieved and characterizes the corresponding scheduling strategies. However, we also uncover scenarios where work-conserving policies fail to attain optimality—an outcome that may be unintuitive to practitioners—thereby revealing the added complexity and subtle trade-offs involved in scheduling for multi-agent LLM systems.
Our contributions include:
  • A formal queuing-theoretic framework for LLM inference scheduling that captures the unique characteristics of the prefill and decode phases, explicitly modeling batching processing time. This modeling effort lowers the barrier for the queuing theory communities to contribute to improving LLM serving systems.
  • A rigorous theoretical analysis establishing that work-conserving scheduling algorithms achieve maximum throughput in single-instance LLM inference, highlighting this as a core design principle for practitioners.
  • An examination of widely used scheduling algorithms reveals that Orca and Sarathi-Serve are throughput-optimal, while FasterTransformer and non-mixed batching in vanilla vLLM can become unstable under moderate load. These findings support the adoption of Sarathi-Serve–type algorithms and offer practical guidance for system selection.
  • An extended analysis on throughput-optimality for AI-agent workloads. Our results reveal key challenges in optimizing throughput in such systems and highlight the need for deeper theoretical and practical investigation in this emerging domain.
Our work emphasizes the practical implications of scheduling algorithm design through both theoretical analysis and empirical validation on real-world workloads. This research lays the foundation for more efficient scheduling algorithm design for LLM serving systems and identifies significant opportunities for cross-disciplinary collaboration between the queuing theory and machine learning systems communities to address the growing demand for LLM inference at scale.
1.1 Background
LLM generative inference. When processing an input request, LLM inference consists of two phases: prefill and decoding. The prefill phase processes the entire request to compute the KV cache (Key-Value cache, which stores the attention keys and values for each layer to avoid redundant computation) and generates the initial response token in a single step. The decoding phase then utilizes the prior contexts or KV cache to generate subsequent tokens one at a time. These phases have distinct resource utilization patterns: prefill is compute-bound, while decoding is memory I/O-bound. Conventional LLM inference engines place both phases on the same GPU group despite their different computational characteristics to maximize resource utilization. While we acknowledge that some recent works propose to disaggregate prefill and decode stages on separate GPUs (Patel et al., 2023; Zhong et al., 2024), the modeling of the p/d-disaggregation is considered analogous to our work.
Inference and serving goal. Two critical metrics evaluate LLM serving performance: throughput and inference latency. Throughput measures the token generation rate within a given timeframe. Inference latency represents the time required to complete a request from start to finish. There are various latency metrics, such as TTFT (time to first token, measuring prefill latency) and TBT (time between token, measuring decoding latency). We evaluate latency performance using service level objective (SLO) attainment, which measures the percentage (e.g., 99%) of requests fulfilled within a predefined timeframe. For online serving, the objective is to optimize for throughput under latency SLO constraints. For offline inference or batch inference, the objective is to optimize batch throughput. In a production cloud system, we observe that from a major cloud service provider that batch or offline inference dominates a major capacity in all workloads, making optimizing for throughput a first order concern.
Inference optimization. To address the challenges of meeting performance requirements while minimizing resource wastes, past work leverage GPU hybrid parallelization strategies, combining data, tensor, and pipeline parallelism to optimize computation across GPUs. Tensor model parallelism (TP)  (Narayanan et al., 2021) distributes computations across GPUs by partitioning transformer weight matrices row-wise and column-wise, with layer outputs aggregated through two AllReduce operations. Pipeline parallelism (PP)  (Huang et al., 2019) segments the model into multiple stages assigned to specific GPUs or GPU groups, with inter-layer activations communicated between stages. Furthermore, advanced memory management techniques, such as paged attention (Kwon et al., 2023), help alleviate memory pressure and improve resource utilization. Scheduler optimizations, including dynamic batching and chunked prefill processing (Agrawal et al., 2023), and disaggregating prompting and decoding phases (Zhong et al., 2024; Patel et al., 2024), ensure high throughput while meeting latency constraints, enabling these systems to efficiently handle diverse LLM inference workloads.
Batching. The computational differences between prefill and decoding phases result in varying performance outcomes when applying batching strategies. The state-of-the-art LLM serving systems currently leverage an optimization technique known as continuous batching (Yu et al., 2022), which combines new request prefill operations with ongoing request decoding operations in the same batch to improve GPU utilization, making batching decisions at the token level. However, this approach creates significant interference between the prefill and decoding phases. When even a single prefill job is introduced into a batch of decoding tokens, it substantially degrades performance for both operations, with the performance penalty becoming more severe as the length of the prefill input increases. There are some works from queuing theory community that investigate the scheduling policy with ML-guided prediction of token length (Mitzenmacher & Shahout, 2025). Our work is complementary and aims to develop a foundation for modeling and scheduling algorithm for LLM and LLM agents.
Queuing theory for batch services. Our work extends the rich tradition of batch queuing theory in several directions. While prior research has explored stochastic service capacity (Chang & Choi, 2005; Janssen & van Leeuwaarden, 2005) and multi-class batch service (Baetens et al., 2018; Reddy et al., 1993), LLM inference presents unique challenges and opportunities: the dual-phase nature of processing with distinct resource profiles, dynamic batch formation across operation types, and interdependencies between request types. By developing a queuing-theoretic framework specifically for LLM inference, we bridge theoretical concepts with practical system design. Our findings on work-conserving policies not only validate certain industry practices but also reveal fundamental limitations in others. Furthermore, our analysis of AI agent workflows extends batch queuing theory into the emerging domain of distributed, collaborative AI systems where traditional assumptions may no longer hold.
2 A stochastic processing model for LLM inference
In this section, we introduce a stochastic processing model that we believe is stingy enough while capturing the realistic scheduling and batching problems for LLM inference. In particular, our model is anchored in Sarathi-serve, a state-of-the-art LLM serving system widely deployed (Agrawal et al., 2023). Yet, it is designed to accommodate general scheduling algorithms, including existing implementations such as those in FastTransformer, vLLM, and Orca (NVIDIA, 2024; Kwon et al., 2023; Yu et al., 2022). We demonstrate how this model facilitates practical insights for designing throughput-optimal algorithms.
2.1 Model Setup
Arrival. In our model, time is assumed to be discrete, indexed by $n\in\mathbb{N}\equiv\{1,2,\ldots,\}$. For time slot $n$, we use $a_{n}$ to denote the number of requests arriving at the system in the time slot: multiple requests arriving at the same time slot are allowed. We envision all requests are ordered. For $i=1,2,\ldots,$ request $i$ is assumed to have $v_{p}(i)\in\mathbb{N}$ (ordered) prefill-tokens, $v_{d}(i)\in\mathbb{N}$ (ordered) decode-tokens, and feature $z(i)\in\mathcal{K}$ where $\mathcal{K}$ is a finite set. Here the feature $z(i)$ can capture the types/contents of a request. $(v_{p}(i),v_{d}(i),z(i))$ is used to describe the system while they may not be observed for the scheduling algorithm (e.g., $v_{d}(i)$ is often not observed). We assume that
$\{a_{n}:n\in\mathbb{N}\}\quad\text{ and }\quad\{(v_{p}(i),v_{d}(i),z(i)),i\in\mathbb{N}\}$ (2.1)
are two iid-sequences, and these two sequences are independent, while $(v_{p}(i),v_{d}(i),z(i))$ is allowed to be correlated. We assume each quantity in (2.1) has the first moment
$\lambda=\mathbb{E}(a_{1}),\quad m_{p}=\mathbb{E}(v_{p}(1)),\quad\text{and}\quad m_{d}=\mathbb{E}(v_{d}(1)).$ (2.2)
Physically, time slots can correspond to the clock frequency unit or other time units that capture the precision of decision-making. [^2] The discrete-time model is introduced for simplicity. Extending it to continuous time is straightforward.
Iteration-level batch serving. We assume the system has one LLM engine that processes both the prefill and decode tokens. The scheduler of LLM engines can select a batch of tokens from multiple requests to process simultaneously for the benefits of computational efficiency. In particular, a certain number of unprocessed tokens from each request can be batched together with the following feasible constraints: (defined mathematically in Eq.(2.5))
  1. (a)
    All prefill tokens in a request must be processed before the first decode token can be loaded for processing. Within both the prefill and decode phases of a request, a token can be loaded into a batch only after the previous token has been processed or has also been included in the batch.
  2. (b)
    No two decode-tokens from the same request can be loaded into a single batch.
Figure 1: Visualization of key scheduling terminologies in LLM engine.
We refer to the total number of loaded tokens $b^{\prime}$ in a batch as token load, and we require $b^{\prime}\leq b$, a predefined token budget. Completing one batch of processing is known as an iteration. The terms iteration, batch, and token budget follow the same definitions as in Sarathi-serve (Agrawal et al., 2023) [^3] In Sarathi-serve, the terms chunk size and token budget are used interchangeably. Furthermore, the number of requests in a batch is referred to as the batch size.. In particular, the authors introduced token budget to ensure iteration times will not have a huge variation from long input prompts to negatively affect TBT of requests in the decode phase.
A scheduling algorithm can select $b^{\prime}$ and determine the token constituents of a batch at each iteration. In fact, most existing LLM serving systems fit within this model. For example, Sarathi-Serve prioritizes decoding requests when loading a batch while allowing prefill tokens to be appended in a first-come, first-served manner up to the token budget $b$ ($b=256$ or $b=512$ is often selected). An example of such a scheduling algorithm is shown in Fig. 1. Other systems (NVIDIA, 2024; Kwon et al., 2023; Yu et al., 2022) differ in their approaches for the batch constituents, such as whether they prioritize prefill or decoding, whether preemption is allowed, and whether mixing batches across decoding and prefill tokens is permitted, among other factors (see Section B.4 for more details).
Figure 2: Batch processing time remains relatively constant for a given token budget (when the LLM is at full token load), and the CoV (coefficient of variation) becomes even smaller given larger models.
Figure 3: Piecewise linear fit for CodeLlama-34B and Llama-70B models for batch processing time under various token budgets and Tensor-parallel sizes (under full load). $R^{2}$ is above 0.985 for all cases.
Batch processing time. A key gap in the current literature is the modeling of batch processing time. While machine learning models exist for processing time prediction (Agrawal et al., 2024), no closed-form approximations are known. It is important to establish closed-form or analytical batch time prediction for downstream scheduling analysis, end-to-end differentiability, etc.
We observe that the processing time of a batch in fact can be well approximated by a piecewise linear function that only depends on the token load $b^{\prime}$
$t_{b^{\prime}}=c+a\cdot\max(0,b^{\prime}-b_{0}).$ (2.3)
Equation (2.3) serves as an empirical formula. Fig. 2 illustrates that $t_{b^{\prime}}$ exhibits minimal variation when batch constituents change while keeping the token load $b^{\prime}$ fixed. Fig. 3 demonstrates how well a linear fit approximates $t_{b^{\prime}}$ as $b^{\prime}$ varies. The parameters $(c,a,b_{0})$ depend on the LLM model and configuration. For example, in Llama-3-70B running on two A100 80GB GPUs with NV Link under tensor parallel, we observe $c=45.5$ ms, $a=0.30$, and $b_{0}=64$ ms.
The strong agreement between equation (2.3) and real processing times is due to the dominance of linear-layer computations (i.e., feed-forward neural layers) in LLM inference (Kamath et al., 2024; Zhu et al., 2024; Ye et al., 2025). In the appendix, we extend this empirical formula to account for attention layer computation and KV-cache memory overhead, to make it even more widely applicable.
We assume batching decisions are made at the start of each time slot, with batch processing spanning multiple slots and completing at the end of a slot. Since batching occurs at the token level, processing is uninterrupted. When the last decode-token of a request finishes at the end of a time slot, the request departs the system.
2.2 A Markov chain
With the model setup complete, we now introduce discrete-time Markov chains (DTMCs) to describe the system’s dynamics. At each time slot $n$, let $\mathcal{Q}_{n}$ denote the set of requests that have not yet departed. We assume that $\{(P_{i}(n),D_{i}(n),Z_{i}(n)):i\in\mathcal{Q}_{n}\}$ is ordered (e.g., by request arrival times) [^4] This assumption is not necessary but allows consideration of scheduling algorithms that utilize order information.. Here, $P_{i}(n)$ and $D_{i}(n)$ represent the number of unprocessed tokens in the prefill and decoding stages, respectively, excluding tokens currently being processed in a batch.
Let $R(n)$ denote the remaining time of the current batch processing. If $R(n)=0$, a new batch can be formed; otherwise, no decision is required. The system state at the beginning of slot $n$ is then defined as
$X(n)=\Big{(}R(n),\{(P_{i}(n),D_{i}(n),Z_{i}(n)):i\in\mathcal{Q}_{n}\}\Big{)}.$ (2.4)
Denote $\mathcal{X}$ as the set of all possible states $X(n)$ can take for each $n\in\mathbb{N}$.
Scheduling algorithms. Let $x=(r,\{(p_{i},d_{i},z_{i}):i\in\mathcal{Q}\})\in\mathcal{X}$ be an arbitrary system state. Given $x$ with $r=0$, a scheduling algorithm selects a batch configuration: $\pi(x)=(\delta_{i}^{p},\delta_{i}^{d})_{i\in\mathcal{Q}}$, where $\delta_{i}^{p}$ and $\delta_{i}^{d}$ denote the number of prefill-tokens and decode-tokens, respectively, from request $i$ to be included in the batch. For $\pi(x)$ to be feasible, it must satisfy:
$p_{i}>0\text{ implies }\delta_{i}^{d}=0$ (2.5a)
$\delta_{i}^{p}\leq p_{i}$ (2.5b)
$\delta_{i}^{d}\leq 1$ (2.5c)
$\sum_{i\in\mathcal{Q}}\Big{(}\delta_{i}^{d}+\delta_{i}^{p}\Big{)}=b^{\prime}\leq b.$ (2.5d)
Eq. 2.5a requires prefill tokens to be completed before decoding, while Eq. 2.5b bounds batched tokens within unprocessed ones. Eq. 2.5c enforces sequential decoding, and Eq. 2.5d sets the token budget constraint. Furthermore, when $r>0$, we enforce $b^{\prime}=0$ since batch processing is ongoing. However, even when $r=0$, an idle time slot (i.e., $b^{\prime}=0$) is allowed. Any $\pi$ satisfying all above conditions is referred to as a scheduling algorithm.
System dynamics. When the system is in state $X(n)=(R(n),\{P_{i}(n),D_{i}(n),Z_{i}(n),i\in\mathcal{Q}_{n}\})$ at time slot $n$, we use $X^{\prime}(n)$ to denote the post-decision state in time slot $n$ before taking into account of the new request-arrivals in time slot $n$. Specially, let $\pi(X(n))=(\delta_{i}^{p},\delta_{i}^{d})_{i\in\mathcal{Q_{n}}}$, we have
$P_{i}^{\prime}(n)=P_{i}(n)-\delta_{i}^{p},i\in\mathcal{Q}_{n},$
$D_{i}^{\prime}(n)=D_{i}(n)-\delta_{i}^{d},i\in\mathcal{Q}_{n},$
$\mathcal{Q}^{\prime}_{n}=\mathcal{Q}_{n}\setminus\{i\in\mathcal{Q}_{n}:D^{\prime}_{i}(n)=0\}.$
Then, the state $X(n+1)$ in time slot $n+1$ after taking into account of the new request-arrivals in time slot $n$ is $X(n+1)=(R(n+1),\{P_{i}(n+1),D_{i}(n+1),Z_{i}(n+1),i\in\mathcal{Q}_{n+1}\}).$ Denoting $\mathcal{A}_{n}$ the set of new request-arrivals in time slot $n$, we have $\mathcal{Q}_{n+1}=\mathcal{Q}^{\prime}_{n}\cup\{\mathcal{A}_{n}\}$ and
$\forall i\in\mathcal{Q}^{\prime}_{n}:\quad\quad P_{i}(n+1)=P_{i}^{\prime}(n),\quad D_{i}(n+1)=D_{i}^{\prime}(n),\quad Z_{i}(n+1)=Z_{i}(n)$
$\forall i\in\mathcal{A}_{n}:\quad\quad P_{i}(n+1)=v_{p}(i),\quad D_{i}(n+1)=v_{d}(i),\quad Z_{i}(n+1)=z^{\prime}(i)$
For the remaining processing time $R(n+1)$, we have
$R(n+1)=\begin{cases}{\color[rgb]{0,0,0}R(n)-1}&R(n)>0\\ {\color[rgb]{0,0,0}t_{b^{\prime}}-1}&R(n)=0\end{cases}$
where $t_{b^{\prime}}$ is the number of slots required for processing the new batch $b^{\prime}:=\sum_{i\in\mathcal{Q}(n)}(\delta_{i}^{p}+\delta_{i}^{d}).$ Assume the iid assumption (2.1), under any scheduling algorithm, $\{X_{n}:n\in\mathbb{N}\}$ is a DTMC.
2.3 Throughput-optimal algorithms
After formalizing the scheduling problem in Secs. 2.1 and 2.2, we are now in a position to analyze the system rigorously. The central question we address is: What is the maximal throughput rate of the system, and which scheduling algorithms can achieve it?
A scheduling algorithm is said to achieve a certain throughput $\lambda$ if it stabilizes the queuing system under that arrival rate in (2.2) is $\lambda$; in other words, the associated discrete-time Markov chain (DTMC) $\{X_{n},n\in\mathbb{N}\}$ is irreducible and positive recurrent when the arrival rate in (2.2) is $\lambda$; see Section 5.7 of (Dai & Harrison, 2020) for a discussion of maximally stable scheduling algorithms. Our main result is that a class of ‘work-conserving’ scheduling algorithms can (almost) achieve the system’s maximal throughput rate.
Specifically, a scheduling algorithm $\pi$ is said to be work-conserving if $\pi(x)$ forms a batch with $b^{\prime}=b$ whenever it is able to. Namely, $\pi(x)$ satisfies
$\sum_{i\in\mathcal{Q}}\Big{(}\delta_{i}^{d}+\delta_{i}^{p}\Big{)}=b$ (2.6)
whenever $\sum_{i\in\mathcal{Q}}\Big{(}p_{i}+1(p_{i}=0)\Big{)}\geq b.$
Under a work-conserving algorithm, whenever possible, $\pi$ will not waste the bandwidth: it will serve up to the token budget. This is achieved by allowing batches that can mix prefill- and decode-tokens.
In proving the stability results for Markov models of a network of LLM engines in Section 4, it is convenient to adopt the fluid limit technique that will be introduced in Section B. Anticipating of using the fluid limit technique, we introduce the following family of scheduling algorithms.
$(K_{p},K_{d})$-FCFS algorithms. Let $K_{p},K_{d}\geq 1$ be integers. When loading tokens in phase $f\in\{p,d\}$, only tokens from (at most) $K_{f}$ oldest requests are loaded, $f\in\{p,d\}$. Among these $K_{f}$ requests, tokens can be selected in any rule including the smallest-remaining-token-first. Having an upper limit of $K_{f}$ requests in phase $f$ is realistic because of the memory constraint in the LLM engine, particularly during the prefill phase. (Our model has completely ignored the memory constraint.) The exact values of $(K_{p},K_{d})$ are not important in our fluid limit analysis.
We now state the theorem formally. In the theorem, we assume each DTMC is irreducible. This condition is easily satisfied by assuming $\mathbb{P}\{a_{n}=0\}>0$ because the work-conserving policy will empty the queue eventually if there is no incoming requests.
Theorem 1.
Assume the iid assumption (2.1) and the first moment assumption (2.2). Assume further the second moment assumption $\mathbb{E}[D_{1}^{2}]<\infty.$ (a) Assume the following system load condition
$\lambda(m_{p}+m_{d})<b/t_{b}.$ (2.7)
Then the DTMC $\{X_{n},n\in\mathbb{N}\}$ is positive recurrent under any work-conserving $(K_{p},K_{d})$-FCFS algorithm. (b) Assume
$\lambda(m_{p}+m_{d})>b/t_{b}.$ (2.8)
Then
$\mathbb{P}\{\lim_{n\to\infty}\lvert X(n)\rvert=\infty\}=1,$ (2.9)
where $\lvert X(n)\rvert=\sum_{i\in\mathcal{Q}_{n}}(P_{i}(n)+D_{i}(n))$ is the total of unprocessed tokens at time slot $n$.
Note that $b/t_{b}$ represents the maximal token processing rate. Therefore, when the system is overloaded—i.e., when condition (B.18) is satisfied—it is evident that no scheduling algorithm can stabilize the system. In practice, under such overload conditions, system memory may become exhausted, leading to requests being blocked or dropped.
While the result in Thm. 1 may appear unsurprising, it highlights two important considerations: (i) In real-world deployments, some systems use scheduling algorithms that are not ‘work-conserving’, and these can be unstable even when the load condition (2.7) is satisfied. We will delve into this further in Sec. B.4, where the theorem offers practical guidance on selecting or modifying scheduling policies to ensure stability. (ii) In Sec. 4, we demonstrate that when LLM instances are connected to a network to handle AI-agent workloads, even work-conserving algorithms may fail to achieve maximal throughput.
These observations underscore that analyzing queue stability is far from trivial—it has long been a central focus in the queuing theory community (Dai & Harrison, 2020; Dai, 1995). Thus, the stability guarantees provided by our theorem not only offer reassurance to handle typical workloads, but also allow practitioners to shift their focus toward optimizing more complex performance metrics such as latency, or tail latency for prefill and decode phases, and addressing challenging system scenarios.
Proof Scketch..
We will provide two proofs for Part a. One proof uses the fluid limit technique that has been standard in the literature (Dai, 1995; Stolyar, 1995; Dai & Harrison, 2020). The fluid limit technique will also be used to prove Part b. The fluid model and fluid limits will be developed in Appendix B.
The second proof works with DTMC directly under the additional assumption that all random variables in (2.1) are bounded. For this second proof of Part a, we construct a Lyapunov function on the system state $X(n)$:
$f(X(n))=\sum_{i\in\mathcal{Q}_{n}}\left(P_{i}(n)+D_{i}(n)\right)+R(n)\cdot\frac{b}{t_{b}}.$
We will show that this Lyapunov function $f$ exhibits a negative drift whenever $X(n)\notin\mathcal{C}$, for some finite set $\mathcal{C}$. By the Foster-Lyapunov criterion, this implies the stability of the system.
A key advantage of this proof technique is that it accommodates a wide class of scheduling algorithms $\pi$, including those that make decisions based on flexible prioritization rules—such as prioritizing requests with the fewest remaining tokens (SJF), or using arbitrary functions of request features and token lengths. Notably, proving stability for shortest-remaining-time-style algorithms has historically been challenging. However, due to the structure of our Lyapunov function and the characteristics of our model, we are able to incorporate such policies into our analysis. Full details are provided in appendix A. ∎
3 Stability issues of incumbent scheduling algorithms
In this section, we formally demonstrate that Sarathi-serve is throughput-optimal, providing reassurance to practitioners. In contrast, FastTransformer and vanilla vLLM are not maximally stable and should be used with caution. We begin by formally describing these systems using the language of our model.
Decode-prioritized schedule (without mixed batching): An example scheduler is FasterTransformer (NVIDIA, 2024). It fills in all the decode tokens from any request $i$ with $p_{i}=0$ until batch size limit with a predefined order (e.g. FCFS, SJF, etc.). No prefill tokens from two different requests are allowed to batch together. This policy should from instability due to the coarse-grained batching (Note that we omit $n$ for brevity).
Prefill-prioritized schedule (without mixed batching): An example scheduler is vLLM (vanilla) (Kwon et al., 2023). It fills in prefill tokens from a request $i$ with $p_{i}>0$ until batch size limit with a predefined order (e.g. FCFS, SJF).
Prefill-prioritized schedule (with mixed batching): An example scheduler is Orca (Yu et al., 2022). It fills in $i$ with $p_{i}>0$ first. Only after $p_{i}=0$ for all $i=1,\ldots,q$, pack as many decode tokens as possible till batch size limit or token budget limit.
Decode-prioritized chunk schedule (Sarathi-Serve): When forming a batch to the processed, fill in as many as decode tokens (at most one decode token for each request) as possible (Agrawal et al., 2023). When there is additional space, fill in as many prompt tokens as possible (from a minimal number of requests). Note that how decode-tokens and prompt-tokens are selected is not important for our results below.
Note: Orca and vLLM both utilize FCFS iteration-level batching with eager admission of prefill requests but differ in how they compose batches. Orca allows hybrid batches that mix prefill and decode requests, whereas vLLM restricts batches to either entirely prefill or entirely decode requests. Despite this distinction, both systems enhance throughput by maximizing batch size until memory capacity bound.
Figure 4: Experimental results demonstrating the instability of FastTransformer and vLLM.
Instability of vanilla vLLM and FastTransformer: From the above formulation, it is evident that Orca and Sarathi both belong to the ‘work-conserving’ class so that they are throughput-optimal in theory (though the practical performance on latency can still vary). Next, we show that FastTransformer and vLLM are unstable even when the load condition (2.7) is satisfied throughout the experiments conducted in Figure 4.
The instability in vLLM arises because it prioritizes prefill over decode. If incoming prefill requests are consistently short, the scheduler becomes non-work-conserving (unable to reach compute bound phase due to short prefills), leading to decode queue buildup.
In contrast, FasterTransformer always prioritizes decode and lacks continuous batching. This means that long decode requests can keep generating new tokens, effectively blocking incoming prefill requests from making progress.
4 Generalization to the AI-agent workload
LLMs are increasingly seen as autonomous agents capable of handling complex tasks and interacting with other agents. In this context, a request may consist of a sequence of LLM calls, possibly involving interactions across different LLM engines. The execution flow of such a request can dynamically change based on LLM outputs. We refer to this type of workload as an AI-agent workload (see Figure 5 for an example). Efficient scheduling and resource optimization for AI-agent workloads are crucial to maximizing throughput and minimizing latency in a world where humans and AI agents coexist and collaborate.
Figure 5: AI-agent workload (Luo et al., 2025)
In Section 2, we define the basic stochastic processing model of the LLM engine. In this section, We show that work-conserving scheduling algorithms can still achieve maximal throughput in several network scenarios. However, inspired by Rybko-Stolyar-type (Rybko & Stolyar, 1992) examples, we also identify settings where work-conserving algorithms may fail, highlighting key challenges in designing AI-agent workflows. We hope this modeling effort encourages operations researchers to contribute to this emerging field.
4.1 Homogeneous, parallel LLM engines
Suppose there are $K$ identical LLM engines. These engines can be configured to work in parallel. Requests can be routed to any one of these LLM engines through a load balancing algorithm; see Figure 7. In this case, the load condition is
$\lambda(m_{p}+m_{d})<Kb/t_{b}.$ (4.1)
One load-balancing algorithm is to assign a request randomly (uniformly) among $K$ engines. Another one is to assign an arriving request to the engine that has the least number of requests. When the information of the numbers of outstanding tokens among all engines is available to the load balance algorithm, one can use the engine that has the least number of tokens.
Proposition 1.
Fix a load balancing algorithm and a work-conserving $(K_{p},K_{d})$-FCFS scheduling algorithm that is used in all engines. The DTMC describing the system is positive recurrent under load condition (4.1).
The proof is in Appendix B.5.
4.2 An LLM engine serving two types of requests
Suppose type $j\in\{1,2\}$ requests has arrival rate $\lambda^{j}$ and mean token sizes $(m^{j}_{p},m^{j}_{d})$. Assume that
$\sum_{j\in\{1,2\}}\lambda^{j}(m^{j}_{p}+m^{j}_{d})<b/t_{p}.$ (4.2)
The fluid model of the basic LLM engine is fully developed in Appendix B. The fluid model for the current LLM engine serving two types of requests can be developed similarly. We claim that the fluid model is stable under any $(K_{p},K_{d})$-FCFS work-conserving scheduling algorithm. Here, the Lyapunov function is still the total workload
$f(t)=\sum_{j}\Big{(}m^{j}_{p}Z^{j}_{p}(t)+(m^{j}_{p}+m^{j}_{d})Z^{j}_{d}(t)\Big{)}.$
Under work-conserving scheduling algorithms, $f(t)>0$ implies that
$\sum_{j}\dot{B}^{j}_{p}(t)+\dot{B}^{j}_{d}(t)=b/t_{b}.$
when $t$ is a regular point. Therefore, $f(t)>0$ implies $\dot{f}(t)=-\delta$, where
$\delta=b/t_{b}-\sum_{j\in\{1,2\}}\lambda^{j}(m^{j}_{p}+m^{j}_{d})>0,$
proving $f(t)=0$ for $t\geq f(0)/\delta$. Therefore, the fluid model is stable as defined in Section B.2. Following the development in Section B.4, the DTMC describing the system is positive recurrent under load condition (4.2).
4.3 A fork-join agent model
Suppose that there are a network of four LLM engines,now called four agents indexed by $j\in\{0,1,2,3\}$. Requests arrive at agent 00 with rate $\lambda$. Upon leaving agent 00, a request forks into two new sub-requests (tasks). Task $1$ will be processed by agent $1$ and task $2$ will be processed by agent $2$. Upon both tasks from the same request are completed, the merged two sub-requests arrives at agent $3$. Assume token sizes at agent $j$ is $(m^{j}_{p},m^{j}_{d})$.
Conjecture 1.
Assume
$\lambda(m^{j}_{p}+m^{j}_{d})<b^{j}/t^{j}_{b}\quad\text{ for each }j\in\{0,1,2,3\}.$ (4.3)
Under any $(K^{j}_{p},K^{j}_{d})$-FCFS work-conserving scheduling algorithm at agent $j$ for $j\in\{0,1,2,3\}$, the DTMC describing the system is positive recurrent.
If we remove agent 3 is the request work flow, which can be “justified” when agent 3 is lightly loaded, it is easy to see the conjecture is true for the resulting $3$-agent system. Indeed, the result for the $3$-agent system can easily be proved using the fluid limit technique in Appendix B.
(a) Total requests comparison
(b) Type A decode tokens comparison
(c) Type A vs Type B in system
(d) Type A prompt tokens comparison
(e) Unstable total requests comparison
(f) Unstable type A decode tokens comparison
(g) Unstable type A vs type B in system
(h) Unstable type A prompt tokens comparison
Figure 6: Comparison between stable and unstable agent behaviors with a work-conserving schedule under loaded condition ($\rho=0.9$).
Assume there are three agents indexed by $j\in\{0,1,2\}$. A request leaving agent 00 has $50\%$ probability triggering a type $A$ sub-tasks and $50\%$ probability triggering a type $B$ sub-tasks A type $A$ request has two sub-tasks, first a small task by agent 1, followed by a longer task by agent 2. After agent 2 completes the second sub-task, the type $A$ request leaves the system. A type $B$ request has two sub-tasks, first a small task by agent 2, followed by a longer task by agent 1. After agent 1 completes the second sub-task, the type $B$ request leaves the system.
Assume that the request arrival rate $\lambda=2$ requests per time slot to agent 00. All agents have token budget $b=10$, and each batch (full or otherwise) takes exactly one time slot to process. Assume that the token sizes are deterministic with $(m^{0}_{p},m_{d}^{0})=(2,2)$ for agent 00,
$(m^{(A,1)}_{p},m^{(A,1)}_{d})=(1,1)\quad\text{ and }\quad(m^{(B,1)}_{p},m^{(B,1)}_{d})=(6,1)$ (4.4)
for agent $1$,
$(m^{(A,2)}_{p},m^{(A,2)}_{d})=(6,1)\quad\text{ and }(m^{(B,2)}_{p},m^{(B,2)}_{d})=(1,1)$ (4.5)
for agent $2$. All agents use work-conserving (mixed batch) scheduling algorithms. For concreteness, agent 00 uses decode-priority FCFS algorithm. Agent $1$ gives non-preemptive priority to type B sub-tasks, Agent $2$ gives non-preemptive priority to type A sub-tasks. Within a type, uses decode-priority FCFS algorithm. For agent $1$, type $A$ and $B$ tokens can be mixed only when type $B$ tokens cannot form a full batch. Similarly, agent $2$ follows a similar rule.
In Figure 6, we show that under the parameters $b=10$, $t_{b}=1$, and token sizes in (4.4) and (4.5) or some variations, the total number of tokens in this system or some variations of the system will go to infinity as time $n\to\infty$. This phenomenon is unique when the processing time has a little noise, but it is not present (a-d) when processing times are deterministic.
5 Latency Optimization
We have developed a class of algorithms that achieve optimal throughput. However, in practice, the algorithm with the lowest latency can vary depending on several factors. While a comprehensive study on latency optimization is left for future work, we present preliminary analyses exploring how the token budget in Sarathi-Serve affects latency performance. These results reveal both challenges and opportunities for further research.
All experiments were conducted using the CodeLlama-34B model. As shown in Fig. 7, both end-to-end (E2E) latency and prefill time are influenced by the choice of token budget in Sarathi-Serve. We find that a moderate token budget (e.g., 512 tokens) yields lower median E2E latency, while a larger token budget (e.g., 1024 tokens) improves prefill latency. However, when the budget is too small (e.g., 128 tokens), overhead increases significantly due to repeated cross-attention across iterations, as indicated by the red line. Figure 8 further demonstrated the reason why under the same policy, certain metrics are better, while other metrics are worse. For example, VLLM has better TBT but worse TTFT.
The key takeaway is that balancing the token budget is essential: larger budgets benefit prefill latency, moderate budgets improve E2E latency, and excessively small budgets lead to significant overhead. In designing scheduling algorithms, it is important to consider not only the token budget but also the distribution of prefill lengths and the service level objectives (SLOs) for the target workloads.
Figure 7: Trade-off between TTFT (Right) and end-to-end latency (Left) with different token budget of the Sarathi scheduling policy, running on CodeLlama-34B with high arrival rate scaled with real production conversation traces.
Figure 8: The request-count dynamics under various scheduling algorithms (Orca, vLLM, and Sarathi). The workload is a scaled version (0.25 the original arrival rate) of the code traces from a production cloud environment on two A100-80G servers with tensor parallelism (Patel et al., 2023).
6 Future Directions and Conclusion
We have focused on throughput and stability of LLM and LLM agents but haven’t comprehensively looked into the optimal scheduling policy under performance measures like tail of TBT or TTFT under various load scenarios (Yu & Scully, 2024). The KV cache memory management for longer context or test-time workloads is also considered future work. Moreover, future work may examine scheduling policies under multi-tenancy with best-effort and latency-critical requests collocated together on the same set of agent models or multi-tenancy models collocated on the same set of physical servers.
Further research into joint optimization techniques that concurrently address model autoscaling, resource allocation with multiple models, KV cache policies, and load-balancing policies may yield substantial improvements in overall system responsiveness and efficiency, particularly in environments characterized by bursty or skewed workload distributions. Collaborative efforts between queuing theorists and system practitioners will be essential to develop models that accurately capture these dynamics and inform the design of next-generation scheduling algorithms.
Appendix A Proof of Theorem 1 Part a
Proof.
Assume $b-\lambda(p+d)=\epsilon^{\prime}.$ Let $\epsilon=\min(\epsilon^{\prime},L\lambda).$ We construct the following Lyapunov function
$f(X)=\sum_{i\in[Q(X)]}\left(p_{i}^{\prime}+d_{i}^{\prime}+\frac{\epsilon(1(p_{i}^{\prime}>0)+d_{i}^{\prime})^{2}}{4L\lambda}\right).$
Let $\pi(X)=(\delta_{i}^{p},\delta_{i}^{d})_{i\in Q(X)}$. Note that $1(p_{i}^{\prime}>0)+d_{i}^{\prime}$ will decrease by 1 only when $\delta_{i}^{p}=p_{i}^{\prime}$ or $\delta_{i}^{d}=1$. Thus the decrease of the Lyapunov function is
$\Delta(X)=\sum_{i\in[Q(X)]}\left((\delta_{i}^{p}+\delta_{i}^{d})+\frac{\epsilon(2(d_{i}^{{}^{\prime}}+1(p_{i}^{\prime}>0))-1)}{4L\lambda}1(\delta_{i}^{p}=p_{i}^{\prime}\text{ or }\delta_{i}^{d}=1)\right)$
Note that for the new arrivals, we have $E[p_{l}+d_{l}]=p+d$ and $E[(d_{l}+1)^{2}]\leq 2E[d_{l}^{2}]\leq 2L$. Hence, the one step change of the Lyapunov function is bounded by
$E[f(X(n+1))]-f(X(n))\leq\lambda(p+d)+\frac{2\epsilon L\lambda}{4L\lambda}-\Delta(X(n)).$
We aim to show that once $f(X(n))\geq B$ for a large enough $B$, we have
$E[f(X(n+1))]-f(X(n))\leq\lambda(p+d)+\frac{\epsilon}{2}-\Delta(X(n))\leq-\frac{\epsilon}{2}.$ (A.1)
If so, we can invoke the classical Foster-Lyapunov Criterion (Theorem 4.3.1 in (Xu, 2023)) to conclude that $X(n)$ is positive recurrent.
To show (A.1) holds, note that if we have $\sum_{i\in[Q(X)]}p_{i}^{\prime}+\sum_{i\in[Q(X)]}1(p_{i}^{\prime}=0)\geq b$, then
$\Delta(X(n))\geq\sum_{i\in[Q(X)]}(\delta_{i}^{p}+\delta_{i}^{d})=b$
by the work-conserving policy. Thus (A.1) holds:
$E[f(X(n+1))]-f(X(n))\leq\lambda(p+d)+\frac{\epsilon}{2}-b\leq\frac{\epsilon}{2}-\epsilon^{\prime}\leq-\frac{\epsilon}{2}.$
Now let’s consider $\sum_{i\in[Q(X)]}p_{i}^{\prime}+\sum_{i\in[Q(X)]}1(p_{i}^{\prime}=0)<b$, and $f(X(n))\geq B>b$. Note that in this setting, every request $i$ in the queue will have either $\delta_{i}^{p}=p_{i}^{\prime}$ (serving the pre-fill part until finishing) or $\delta_{i}^{d}=1$ (serving the decoding part by $1$). Otherwise the work-conserving assumption is violated.
We establish the proof by the following intuition: when $f(X(n))$ is large enough, there exists a request with sufficiently large decoding length $\tilde{d}$. This request will be served and the decrease of the Lyapunov function $\Delta(X)$ will be lower bounded by $\frac{\epsilon(2\tilde{d}-1)}{4L\lambda}\geq b$, which will be sufficient for showing (A.1).
In particular, let $B:=2b\left(\frac{4L\lambda}{\epsilon}b\right)^{2}+b$ and $\tilde{d}:=\max_{i\in[Q(X)]}d_{i}^{\prime}.$ We have
$f(X(n))\geq B$ $\implies\sum_{i\in[Q(X)]}\left(p_{i}^{\prime}+d_{i}^{\prime}+\frac{\epsilon(d_{i}^{\prime}+1)^{2}}{4L\lambda}\right)\geq B$
$\overset{(i)}{\implies}\sum_{i\in[Q(X)]}d_{i}^{\prime}+\frac{\epsilon(d_{i}^{\prime})^{2}}{L\lambda}\geq B-b$
$\overset{(ii)}{\implies}\tilde{d}+\frac{\epsilon\tilde{d}^{2}}{L\lambda}\geq\frac{(B-b)}{b}$
$\overset{(iii)}{\implies}2\tilde{d}^{2}\geq\frac{(B-b)}{b}$
$\implies\tilde{d}\geq\frac{4L\lambda}{\epsilon}b$
where (i) is due to that $\sum_{i\in[Q(X)]}p_{i}^{\prime}<b$ and $d_{i}^{\prime}+1\leq 2d_{i}^{\prime}$ for $d_{i}^{\prime}\geq 1$; (ii) is due to that $\sum_{i\in[Q(X)]}p_{i}^{\prime}+\sum_{i\in[Q(X)]}1(p_{i}^{\prime}=0)<b$ implies $Q(X)<b$ (the number of requests is bounded by $b$ since each request can be served by at least one token) and the pigeonhole principle; (iii) is due to that $d_{i}^{\prime}\leq d_{i}^{\prime 2}$ and $\epsilon\leq L\lambda$ by the definition of $\epsilon$. This provides a lower bound for $\tilde{d}$.
Next, note that $\tilde{d}$ also provide a lower bound for $\Delta(X)$ because all requests in the queue satisfy $\delta_{i}^{p}=p_{i}^{\prime}\text{ or }\delta_{i}^{d}=1$:
$\Delta(X(n))\geq\frac{\epsilon(2\tilde{d}-1)}{4L\lambda}\geq\frac{\epsilon\tilde{d}}{4L\lambda}.$
Then when $f(X(n))\geq B$, we have
$\Delta(X(n))\geq\frac{\epsilon\tilde{d}}{4L\lambda}\geq\frac{\epsilon}{4L\lambda}\frac{4L\lambda}{\epsilon}b=b.$
Thus (A.1) holds and we complete the proof. ∎
Appendix B The fluid model of the LLM engine and Proof of Theorem 1
Recall the DTMC model of the LLM engine introduced in Section 2. It is convenient to repeat the load condition (2.7) here, namely
$\lambda(m_{p}+m_{b})<b/t_{p},$ (B.1)
where $b$ is token budget, $m_{p}$ and $m_{b}$ are mean prefill and decode token sizes of a request, $t_{p}$ is the processing time of a full batch, and $\lambda$ is the request arrival rate. In this appendix, we will first introduce the dynamics of the processing model of the LLM engine in Section B.1. We will next define the fluid model of the LLM engine and prove the fluid model operating under any work-conserving scheduling algorithm is stable under load condition (B.1) in Section B.2. In Section B.3, we introduce fluid limits that justify the fluid model equations in the fluid model. In Section B.4, we give a proof of Theorem 1 in Section 2.
Although Theorem 1 has been proved in Appendix A without invoking fluid limit technique when random variables in (2.1) are bounded. However, in the setting of AI agents served by a network of LLM engines, fluid limit technique will be useful, even critical, to study maximal stability of many scheduling algorithms. Readers are referred to (Dai & Harrison, 2020) and references there for a detailed coverage of fluid limit technique for studying stochastic processing network stability.
B.1 System dynamics and scheduling algorithms
To study the dynamics of the processing model of the LLM engine, we keep track of the following quantities. Let $Z_{p}n)$ and $Z_{d}(n)$ be the number of requests in phase $p$ and phase $d$, respectively, at the end of time slot $n$, after accounting for the new arrivals and the service completions in time slot $n$. Let $W_{p}(n)$ and $W_{d}(n)$ be the number of prefill tokens in phase $p$ and decode tokens in phase $d$ at time $n$, including those in a batch being processed.
Assume a batch is completed at $n$ and is ready to load the next batch to be started at time $n+1$. The LLM engine needs a scheduling algorithm to decide which tokens to go into the batch.
The following scheduling algorithms can be used to load the next batch. Using newly introduced notation, we recap the definitions of two families of scheduling algorithms.
Work-conserving scheduling algorithms. When
$W_{p}(n)+Z_{d}(n)\geq b,$ (B.2)
the next batch is a full batch.
To describe the dynamics, we further introduce the quantities $B_{f}(n)$, $E(n)$, $F_{f}(n)$, and $V_{f}(n)$. For each phase $f\in\{p,d\}$, define $B_{f}(n)$ to be the cumulative number of tokens that have been completed phase $f$ processing by time $n$; define $F_{f}(n)$ to be the cumulative number of requests that have completed phase $f$ processing by time $n$; define $E(n)=\sum_{\ell=1}^{n}a_{\ell}$ to be the cumulative number of requests that have arrived by $n$, and
$V_{f}(N)=\sum_{i=1}^{N}v_{f}(i)$ (B.3)
to be the total number of phase $f$ tokens brought in by the first $N$ requests. It follows that
$Z_{p}(n)=E(n)-F_{p}(n)\geq 0,$ (B.4)
$Z_{d}(n)=F_{p}(n)-F_{d}(n)\geq 0,$ (B.5)
$W_{p}(n)=V_{p}(E(n))-B_{p}(n)\geq 0,$ (B.6)
$W_{d}(n)=V_{d}(F_{p}(n))-B_{d}(n)\geq 0,$ (B.7)
$B_{f}\big{(}n_{2}t_{b}\big{)}-B_{f}\big{(}n_{1}t_{b}\big{)}\leq(n_{2}-n_{1})b,$
[…]
Under any $(K_{p},K_{d})$-FCFS scheduling algorithm,
$V_{f}\big{(}F_{f}(n)-K_{f}+1\big{)}$ $-(K_{f}-1)\max_{1\leq i\leq E(n)}v_{f}(i)$
$\leq B_{f}(n)\leq V_{f}\big{(}F_{f}(n)+K_{f}\big{)},\quad f\in\{p,e\}.$ (B.9)
Relationship (B.1) is identical to the Key Relationship (6.51) in (Dai & Harrison, 2020)., which plays an important role in defining fluid limits.
B.2 The LLM fluid model & fluid model calculus
In this section, we introduce the fluid model of the LLM engine. The fluid model is defined through a set of fluid model equations including some inequalities. These equations will need to be justified through a fluid limit procedure to be explained in Section B.3. Fix any $(K_{p},K_{d})$-FCFS scheduling algorithm. These fluid model equations include: for any time $t\in\mathbb{R}_{+}\equiv[0,\infty)$,
$Z_{p}(t)=Z_{p}(0)+\lambda t-F_{p}(t),$ (B.10)
$Z_{d}(t)=Z_{d}(0)+F_{p}(t)-F_{d}(t),$ (B.11)
$B_{f}(0)=0,\qquad 0\leq B_{f}(t)-B_{f}(s)\leq(t-s)b/t_{b},\quad 0\leq s\leq t,\quad f\in\{p,d\}.$ (B.12)
$F_{f}(t)=\frac{1}{m_{f}}B_{f}(f),\quad f\in\{p,d\},$ (B.13)
$W_{f}(t)=m_{f}Z_{f}(t),\quad f\in\{p,d\}.$ (B.14)
Here, we intentionally overload the notational system in Section B.1 to emphasize the similarity of the fluid analog. Each overloaded function is differentiated by its argument $n\in\mathbb{N}$ and $t\in\mathbb{R}_{+}$. Fluid model equation (B.12) says that the function $B_{f}(\cdot)$ is Lipschitz continuous with Lipschitz constant $b/t_{b}$. Fluid model equation (B.13) implies that function $F_{f}(\cdot)$ is Lipschitz continuous. Fluid model equations (B.10)-(B.11) and (B.14) imply that $Z_{f}(\cdot)$ and $W_{f}(\cdot)$ are Lipschitz continuous as well. In conclusion, the multidimensional function
$\big{(}B_{f}(\cdot),F_{f}(\cdot),W_{f}(\cdot),Z_{f}(\cdot),f\in\{p,d\}\big{)}$ (B.15)
is Lipschitz continuous. Therefore, from real analysis (c.f. Lemma A.2 of (Dai & Harrison, 2020). and its commentary), this function is absolutely continuous. As a consequence, it is differential almost surely everywhere, and for any component $x(\cdot)$ in (B.15),
[…]
where $\dot{x}(u)$ denotes the derivative of the function $x(\cdot)$ at time $u$. Eequation (B.16) is simply the fundamental theorem of calculus when $x(\cdot)$ is continuously differential. In general, the integral on the right of (B.16) is interpreted as the Lebesgue integral, and it is sufficient that $\dot{x}(u)$ is well defined almost everywhere. See, for example, Lemma A.3 of (Dai & Harrison, 2020). and the references there.
Definition 1.
A point $t>0$ is said to be a regular point for a fluid model solution (B.15) if the solution is differential at time $t$.
When the fluid model solution is clear in a context, we simply say a time $t$ a regular point without reference to the fluid model solution. When the $(K_{p},F_{d})$-FCFS scheduling algorithm is also work-conserving, the fluid model equations include: for any regular time $t>0$,
$W_{p}(t)+W_{d}(t)>0\text{ implies }\dot{B}_{p}(t)+\dot{B}_{d}(t)=b/t_{b}.$ (B.17)
Definition 2.
A function in (B.15) is said to be a fluid model solution if it satisfies fluid model equation (B.10)-(B.14), a work-conserving fluid model solution if, in addition, it satisfies (B.17).
Definition 3.
The fluid model is said to be stable if there exists a time $\delta>0$ such that $Z(t)=0$ for $t\geq\delta$ for any fluid model solution with $\lvert Z(0)\rvert\equiv Z_{p}(0)+Z_{d}(0)\leq 1$.
Definition 4.
The fluid model is said to be weakly unstable if there exists a time $a>0$ such that for each fluid model solution starting from 00, $Z_{p}(a)+Z_{d}(a)\not=0$.
Proposition 2.
(a) Under any work-serving, $(K_{p},K_{d})$ FCFS algorithm, the fluid model is stable under load condition (B.1). (b) Assume
$\lambda(m_{p}+m_{d})>b/t_{p}.$ (B.18)
Under any scheduling algorithm, the fluid model is weakly unstable.
Proof.
Given any fluid model solution $(B(\cdot),F(\cdot),W(\cdot),Z(\cdot))$, we use Lyapunov function
$f(t)=m_{p}Z_{p}(t)+(m_{p}+m_{d})Z_{d}(t)$ (B.19)
as the total workload in the system at time $t$. One interpretation of the total workload is the time needed for the LLM engine to clear off all existing “fluid tokens” in the system when the request arrival is turned off. It follows from fluid model equations (B.10)-(B.14) that
$f(t)=f(0)+\lambda(m_{p}+m_{d})t-\big{(}B_{p}(t)+B_{d}(t)\big{)}.$
Proof of (a). For each regular point $t>0$, $f(t)>0$ implies that $\dot{f}(t)=\lambda(m_{p}+m_{d})-b/t_{p}$, which is negative by load condition (B.1). It follows from Lemma 8.5 of (Dai & Harrison, 2020). that $f(t)=0$ for $t\geq f(0)/\delta$, where
$\delta=b/t_{p}-\lambda(m_{p}+m_{d})>0.$
Proof of (b). Let $t>0$ be a regular point. Because
$\dot{B}_{p}(t)+\dot{B}_{d}(t)\leq b/t_{b},$
$\dot{f}(t)\geq\lambda(m_{p}+m_{d})-b/t_{p}>0$. Since $f(0)$ is assumed to zero, it follows that
[…]
Thus, the fluid model is weakly unstable by choosing any $a>0$. ∎
B.3 Fluid limit
Fix a work-conserving $(K_{p},K_{p})$-FCFS scheduling algorithm. In this section, we define fluid limits and prove that each fluid limit is a fluid model solution satisfying (B.10)-(B.17). For each state $x$ in the space $\mathcal{X}$ in Section 2.2, we use $z_{p}$ and $z_{d}$ to denote the corresponding token counts in prefill phase and decode phase, respectively. Define $\lvert x\rvert=z_{p}+z_{d}$ to be total number of tokens in state $x$. Assume that two iid sequences in (2.1) are defined on some probability space $(\Omega,\mathcal{F},\mathbb{P})$. We make a slightly stronger moment assumption on the token size distribution. There exists an $\epsilon>0$ such that
$\mathbb{E}\Big{[}(v_{f}(1))^{1+\epsilon}\Big{]}<\infty\quad f\in\{p,d\}.$ (B.20)
Clearly, both light and heavy tail token sizes are allowed. Under condition (B.20), Proposition B.8 implies the following
$\mathbb{P}\Big{\{}\lim_{N\to\infty}\frac{1}{N}\max_{1\leq i\leq N}\big{(}v_{p}(i)+v_{d}(i)\big{)}=0\Big{\}}=1.$ (B.21)
Recall the definitions in Section B of $B_{f}(n)$, $E(n)$, $F_{f}(n)$, $W_{f}(n)$, and $Z_{f}(n)$ for each $n\in\mathbb{N}$ and each phase $f\in\{p,d\}$. For each sample path $\omega\in\Omega$, one has realization of two sequences $\{a_{n}(\omega),n\in\mathbb{N}\}$ and $\{(v_{p}(i,\omega),v_{d}(i,\omega)),i\in\mathbb{N}\}$. With these two sequences, the given initial state $x\in\mathcal{X}$, and any fixed $K$-FCFS scheduling algorithm, one can construct the corresponding realization $B^{x}_{f}(n,\omega)$, $E(n,\omega)$, $F^{x}_{f}(n,\omega)$, $W^{x}_{f}(n,\omega)$, and $Z^{x}_{f}(n,\omega)$ for each $n\in\mathbb{N}$ and each $f\in\{p,d\}$. For each $\omega\in\Omega$, each time $t\geq 0$, and each state $x\in\mathcal{X}$ with $\lvert x\rvert>0$, define fluid scaled quantities
$\hat{E}^{x}(t,\omega)=\frac{1}{\lvert x\rvert}E(\lvert x\rvert t,\omega),\quad\hat{B}^{x}_{f}(t,\omega)=\frac{1}{\lvert x\rvert}B^{x}_{f}(\lvert x\rvert t,\omega),\quad\hat{F}^{x}_{f}(t,\omega)=\frac{1}{\lvert x\rvert}F^{x}_{f}(\lvert x\rvert t,\omega),$
$\hat{W}^{x}_{f}(t,\omega)=\frac{1}{\lvert x\rvert}W^{x}_{f}(\lvert x\rvert t,\omega),\quad\hat{Z}^{x}_{f}(t,\omega)=\frac{1}{\lvert x\rvert}Z^{x}_{f}(\lvert x\rvert t,\omega),$
where, whenever $\lvert x\rvert t$ is not an integer, it is applied the floor operation to make it an integer. Following the SLLN theorem,
$\mathbb{P}\Big{\{}\omega\in\Omega:\lim_{n\to\infty}\frac{1}{n}E(n,\omega)=\lambda,\quad\lim_{N\to\infty}\frac{1}{N}V_{p}(N,\omega)=m_{p},\quad\lim_{N\to\infty}\frac{1}{N}V_{d}(N,\omega)=m_{d}\Big{\}}=1.$ (B.22)
Denote the set $\Omega_{0}\subset\Omega$ in which the limits in (B.22) and (B.21) exist. Clearly $\mathbb{P}\{\Omega_{0}\}=1$. We now define fluid limits for each sample path $\omega\in\Omega_{0}$.
Lemma 1.
For each $\omega\in\Omega_{0}$ and each unbounded set $A\subseteq\mathcal{X}$, there exists a sequence $\{x_{k}\}\subset A$ with $\lvert x_{k}\rvert\to\infty$ as $k\to\infty$, and a function $(\hat{B}_{p}(\cdot),\hat{B}_{d}(\cdot))$ such that
$\lim_{k\to\infty}\sup_{0\leq t\leq T}\lvert B^{(x_{k})}_{f}(t,\omega)-\hat{B}_{f}(t)\rvert=0\quad\text{for each }T>0\text{ and }f\in\{p,d\}.$ (B.23)
The proof follows from the Lipschitz property (B.8) and is identical to the proof of (12.37) on page 285 of (Dai & Harrison, 2020). The convergence mode in (B.23) is known as the uniform convergence on compact sets or u.o.c. convergence. Since $\lvert\hat{Z}^{(x_{k})}(0,\omega)\rvert=1$, the sequence $\{\hat{Z}^{(x_{k})}(0,\omega),k\geq 1\}$ is in a compact set of +$\mathbb{R}^{2}_{+}$. By taking a subsequence if needed, we will assume
$Z^{(x_{k})}(0,\omega)\to z=(z_{p},z_{d})$ (B.24)
for some $z\in\mathbb{R}^{2}_{+}$ for the sequence of $x_{k}$ in (B.23).
Lemma 2.
Under any $(K_{p},K_{d})$-FCFS scheduling algorithm, fix a $\omega\in\Omega_{0}$ and a sequence $\{x_{k}\}\subset\mathcal{X}$ such that (B.23) and (B.24) hold. For each $f\in\{p,d\}$,
$\Big{(}\hat{F}^{x_{k}}_{f}(\cdot,\omega),\hat{W}^{x_{k}}_{f}(\cdot,\omega),\hat{Z}^{x_{k}}_{f}(\cdot,\omega)\Big{)},\to\Big{(}\hat{F}_{f}(\cdot),\hat{W}_{f}(\cdot),\hat{Z}_{f}(\cdot)\Big{)}\quad u.o.c.\ \quad\text{ as }k\to\infty,$ (B.25)
where
$\hat{F}_{f}(t)=\frac{1}{m_{f}}\hat{B}_{f}(t),$ (B.26)
$\hat{Z}_{p}(t)=z_{p}+\lambda t-\hat{F}_{p}(t),$ (B.27)
$\hat{Z}_{d}(t)=z_{d}+\hat{F}_{p}(t)-\hat{F}_{d}(t),$ (B.28)
$\hat{W}_{f}(t)=m_{f}\hat{Z}_{f}(t).$ (B.29)
Proof.
To prove (B.25), it is sufficient to prove the u.o.c. convergence for each component. For $\hat{F}_{f}^{x_{k}}(\cdot,\omega)$, we utilize relationship (B.1). The proof is given by Lemma 6.8 of (Dai & Harrison, 2020). and fluid model equation (B.26) is identical to (6.43) of (Dai & Harrison, 2020). The rest of the proof is easy to complete, following (B.4)-(B.7) and analogous to the proof of Theorem 12.13 in (Dai & Harrison, 2020). ∎
Definition 5.
Functions $(\hat{B}_{f}(\cdot),\hat{F}_{f}(\cdot),W_{f}(\cdot),Z_{d}(\cdot))$ is called a fluid limit if there exists $\omega\in\Omega_{0}$ and a sequence of initial states $\{x_{n}\}\subset\mathcal{X}$ such that $\lvert x\rvert_{k}\to\infty$ and limits (B.23) and (B.25) hold.
Lemma 3.
Under any $(K_{p},K_{d})$-FCFS work-conserving scheduling algorithm, each fluid limit satisfies fluid model equations (B.10)-(B.17).
Proof.
The fluid limit satisfies fluid model equations (B.10)-(B.15) is covered in (B.26)-(B.29). It suffices to prove that each work-conserving fluid limit satisfies (B.17). The proof of the latter is analogous to the proof of Theorem 7.2 of (Dai & Harrison, 2020). ∎
B.4 Proof of Theorem 1
Definition 6.
Under a fixed scheduling algorithm, the fluid limits are said to be stable if there exists a constant $\delta>0$ such that
$\hat{Z}_{f}(t)=0\quad t\geq\delta$ (B.30)
for each fluid limit $(\hat{B},\hat{F},\hat{W},\hat{Z})$ defined in Definition 5.
Proposition 3.
Fix a $(K_{p},K_{d})$-FCFS work-conserving scheduling algorithm. If the fluid limits are stable, then the corresponding DTMC in Section 2.2 is positive recurrent.
Proof.
The proof is analogous to the proof of Theorem 6.2 of (Dai & Harrison, 2020). The latter theorem is stated in the setting of continuous time Markov chains. Assume the fluid limits are stable. The Lyapunov function used in Lemma 3.7 of (Dai & Harrison, 2020). in the continuous setting can be copied verbatim in the current setting; see the proof of Theorem 12.27 in (Dai & Harrison, 2020)., which is for the slotted discrete time model. In the proof of Theorem 6.2 in (Dai & Harrison, 2020)., one needs to verify that the sequence of random variables
$\left\{\frac{E(n)}{n},n\geq 1\right\}\text{ is uniformly integerable.}$. (B.31)
In (Dai & Harrison, 2020)., Proposition B.5 is cited to prove (B.31). The proof of Proposition B.5 there utilized the finiteness of the second moment of $a_{n}$ in (2.1). Examining the proof of Lemma 4.5 of (Dai, 1995), one concludes that Proposition B.5 continues to hold under the first moment assumption on $a_{n}$ as in (2.2). ∎
Proof of Theorem 1..
Part (a). By Lemma 3, each fluid limit $(\hat{B},\hat{F},\hat{W},\hat{Z})$ satisfies $\lvert Z(0)\rvert=1$ and fluid model equations (B.10)-(B.17). The fluid model has been shown stable under the load condition (B.1) in Part (a) of Proposition 2. Therefore the fluid limits are stable. By Proposition 3, the DTMC is positive recurrent.
Part (b). Assume condition (2.8). Part (b) of Proposition 2 shows the the fluid model is weakly unstable. Lemma 3 shows that each fluid limit is a fluid solution. Therefore, the proof of Theorem 3.2 of (Dai & Weiss, 1996) can be adapted to conclude Part (b) of the theorem. ∎
B.5 A Sketch of the Proof of Proposition 1
Proof.
The fluid limit technique for proving Theorem 1 continues to apply here. Let $(Z^{k}_{p}(t),Z_{d}^{k}(t))$ be the fluid request level at server $k$. Let $A_{k}(t)$ be the cumulative amount of fluid request to server $k$ by time $t$. Define
$f^{k}(t)=m_{p}Z^{k}_{p}(t)+(m_{p}+m_{d})Z^{k}_{d}(t).$
Under the random assignment load-balancing algorithm, at each regular point, $\dot{A}^{k}(t)=\lambda/K$. Assume $f^{k}(t)>0$,
$\dot{f}^{k}(t)=(m_{p}+m_{d})\lambda/K-b/t_{p}<0.$
Thus $f^{k}(t)=0$ for $t\geq f^{k}(0)/\delta$ and $k=1,\ldots,K$, where $b/t_{b}-(m_{p}+m_{d})\lambda/K>0$. It follows that $Z(t)=0$ for $t>\max(f^{k}(0))/\delta$, proving the fluid model is stable.
Under the join-lowest-request load balancing algorithm, define
$f(t)=\max_{k}f^{k}(t).$
Suppose that $f^{\ell}(t)>\max_{k\neq\ell}f^{k}(t)$. At each regular point of $t$ that is also a differential point of $f$,
$\dot{f}(t)=\dot{f}^{\ell}(t)$
Since $\dot{A}^{\ell}(t)=0$,
$\dot{f}^{\ell}(t)=-b/t_{p}.$
Therefore, $\dot{f}(t)=-b/t_{p}$.
Suppose that $f^{\ell_{1}}(t)=f^{\ell_{2}}(t)>\max_{k\not\in\{\ell_{1},\ell_{2}\}}f^{k}(t)$. One can also prove $\dot{f}(t)=\dot{f}^{\ell_{1}}(t)=-b/t_{p}$.
Suppose that $f^{1}(t)=\ldots=f^{K}(t)$. In this case,
$\dot{f}(t)=\dot{f}^{k}(t)\text{ and }f^{k}(t)>0,\quad k=1,\ldots,K$
Thus,
$\dot{f}(t)$ $=\frac{1}{K}\sum_{k}\dot{f}^{k}(t)=\frac{1}{K}\sum_{k}\Big{(}\dot{A}^{k}(t)(m_{p}+m_{d})-b/t_{p}\Big{)}$
$=\frac{1}{K}(m_{p}+m_{d})\sum_{k}\dot{A}^{k}(t)-b/t_{p}.$
Because $\sum_{k}\dot{A}^{k}(t)=\lambda$, one has $\dot{f}(t)=-\delta<0$. ∎